# MoRE-Brain: Routed Mixture of Experts for Interpretable and Generalizable Cross-Subject fMRI Visual Decoding

### v2

@@ Abstract
Decoding visual experiences from fMRI offers a powerful avenue to understand human perception and develop advanced brain-computer interfaces. However, current progress often prioritizes maximizing reconstruction fidelity while overlooking interpretability, an essential aspect for deriving neuroscientific insight. To address this gap, we propose MoRE-Brain, a neuro-inspired framework designed for high-fidelity, adaptable, and interpretable visual reconstruction. MoRE-Brain uniquely employs a hierarchical Mixture-of-Experts architecture where distinct experts process fMRI signals from functionally related voxel groups, mimicking specialized brain networks. The experts are first trained to encode fMRI into the frozen CLIP space. A finetuned diffusion model then synthesizes images, guided by expert outputs through a novel dual-stage routing mechanism that dynamically weighs expert contributions across the diffusion process. MoRE-Brain offers three main advancements: First, it introduces a novel Mixture-of-Experts architecture grounded in brain network principles for neuro-decoding. Second, it achieves efficient cross-subject generalization by sharing core expert networks while adapting only subject-specific routers. Third, it provides enhanced mechanistic insight, as the explicit routing reveals precisely how different modeled brain regions shape the semantic and spatial attributes of the reconstructed image. Extensive experiments validate MoRE-Brain’s high reconstruction fidelity, with bottleneck analyses further demonstrating its effective utilization of fMRI signals, distinguishing genuine neural decoding from over-reliance on generative priors. Consequently, MoRE-Brain marks a substantial advance towards more generalizable and interpretable fMRI-based visual decoding. Code will be publicly available soon: https://github.com/yuxiangwei0808/MoRE-Brain.
@@ 1 Introduction
Understanding brain visual systems and human perception has been a central interest in the field of neuroscience [1, 2]. The human visual system transforms perceived scenes into rich, meaningful perception through a complex cascade of hierarchical and specialized processing [3]. With the advancement in the diffusion models [4] and representational space like CLIP [5], numerous studies propose to decode visual stimuli from non-invasive brain signals such as functional MRI (fMRI) and EEG, aiming to unravel the complex mechanism under perception and advance the practical brain-computer interface. These methods usually leverage strong priors from existing generative models and train dedicated decoders to decode brain signals to frozen representational space, reconstructing natural scenes with impressive structural accuracy [6, 7, 8, 9, 10, 11].
@@
Despite this progress, a significant gap persists between technical feats and neuroscientific insight. Current leading methods often treat fMRI signals as relatively monolithic inputs, lacking architectural designs that reflect established principles of neural processing. For instance, several works manually segment the brain into broad regions like the higher visual and lower visual cortex, then decode semantic and low-level spatial information from them separately [12, 13, 14]. However, this simplification overlooks the intricate nature of the visual pathway, which involves multiple hierarchical processing stages across numerous specialized cortical areas, each handling different facets of visual information [15, 16]. Furthermore, leading research commonly encounters two critical bottlenecks: cross-subject generalization and limited mechanistic interpretability. Due to inter-individual brain variability, models often require subject-specific training. While approaches like mapping subject data to a shared space [6, 17, 11] or utilizing subject-specific tokens [9] exist, they typically necessitate extra training and are difficult to generalize to new subjects. More fundamentally, the predominant focus on maximizing decoding performance often obscures how different neural computations contribute to the final reconstruction. This lack of transparency hinders our ability to validate these models against neuroscientific knowledge or use them to deepen our understanding of the visual system itself.
@@
To bridge the gap, we introduce MoRE-Brain (Mixture-of-Experts Routed Brain decoder), a novel visual decoding framework that is explicitly designed to mirror the hierarchical and specialized nature of the human visual pathway [18] to overcome key limitations in generalization and interpretability inherent in less biologically constrained approaches. MoRE-Brain builds upon a hierarchical Mixture-of-Experts (MoE) architecture to first map fMRI activity patterns to image and text priors. Instead of relying on a single, monolithic network, MoRE-Brain employs multiple layers of "expert" subnetworks. Within each layer, trainable "router" networks learn to assign specific brain regions (voxel groups) to the experts best suited to process their signals. This allows each expert to specialize in decoding information from distinct neural populations, analogous to functional specialization in the cortex. Addressing the challenge of cross-subject generalization, we draw from findings suggesting that significant inter-individual variation arises from spatial differences in functional network topography rather than fundamental computational differences [19, 20]. Based on this, we hypothesize that the core visual decoding computations (the experts) can be shared across subjects, while subject-specific adjustments are primarily needed for the routers that map individual brain topographies to these experts. Our experiments validate this, demonstrating that by freezing expert weights and fine-tuning only the routers, MoRE-Brain achieves strong cross-subject performance with minimal subject-specific data, offering a more scalable approach than retraining entire models.
@@
Furthermore, effectively utilizing the disentangled information captured by these specialized experts within the iterative image generation process of diffusion models presents another challenge. This mechanism operates dynamically across both time and space: (1) A Time Router selects which hierarchical level of expert embeddings to prioritize at different diffusion timesteps. This potentially reflects the coarse-to-fine dynamics of visual processing, where global scene layout might be established before finer details emerge [21]. (2) A Space Router then dynamically modulates the spatial influence of the outputs from individual experts within the selected level. This mimics how the brain might integrate features processed by different specialized areas (e.g., form, color, motion) into a spatially coherent percept [22]. Critically, the dual-routing system not only guides image generation but also provides a mechanistic lens into the reconstructing process. By observing which experts and levels are routed at different times and for images with different semantic concepts, we gain insight into how distinct, modeled neural sources dynamically contribute to the final reconstruction.
@@ 2.1 Preliminaries: Latent Diffusion Models and Conditioning
Our reconstruction module leverages latent diffusion models (LDMs), specifically Stable Diffusion XL (SDXL) [4], which are generative models operating in a compressed latent space [23, 24]. LDMs learn to reverse a diffusion process that gradually adds Gaussian noise to latent representations ($z_{0}$) over discrete timesteps $t=1,\dots,T$. The forward process is defined as $z_{t}=\sqrt{\bar{\alpha}_{t}}z_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$, where $\bar{\alpha}_{t}$ is a noise schedule and $\epsilon\sim\mathcal{N}(0,\mathbf{I})$. The core of the LDM is a denoising network $\epsilon_{\theta}$ (typically a U-Net [25]), trained to predict the noise $\epsilon$ added at timestep $t$, given the noisy latent $z_{t}$ and conditioning information $c$:
$\mathcal{L}_{LDM}=\mathbb{E}_{z_{0},\epsilon,c,t}\left[||\epsilon-\epsilon_{\theta}(z_{t},t,c)||^{2}_{2}\right]$ (1)
Conditioning information $c$ is typically derived from various modalities. For text-to-image synthesis, $c$ often consists of embeddings $\tau_{text}(c_{text})$ from a pre-trained text encoder (e.g., CLIP [5]). Image conditioning can be achieved using IP-Adapter [26], where features $\tau_{img}(c_{img})$ are extracted from a reference image via a pre-trained image encoder (e.g., CLIP image encoder) and mapped through an adapter network. Both $\tau_{text}(c_{text})$ and $\tau_{img}(c_{img})$ are integrated into $\epsilon_{\theta}$ usually via cross-attention mechanisms. In MoRE-Brain, we adapt this conditioning mechanism to use fMRI-derived embeddings as image and text priors.
@@ 2.2 MoRE-Brain Framework Overview
  • Dynamically Conditioned Image Generation: We introduce a novel Time and Space routing mechanism to dynamically select and integrate the multi-level expert embeddings produced in Stage 1 to guide the diffusion model’s (SDXL’s) denoising process at each timestep $t$. During this stage, the MoE fMRI encoder (trained in Stage 1) is frozen, while the SDXL U-Net (via LoRA [27]) and the Time and Space routers are fine-tuned.
@@ 2.3 Hierarchical MoE for fMRI Encoding
For progressively finer specialization, outputs $O_{j}^{(l)}$ from experts at level $l$ serve as input $X^{(l+1)}$ to the routers at the next level $(l+1)$. We start with $e_{0}=2$ experts at the first level, and each expert’s output path is routed to 2 new experts at the subsequent level. We use a total of $L=4$ levels, resulting in $162^{4}=$ 16 experts at the final level. This number was empirically chosen to approximate the number of distinct functional ROIs identified in recent visual cortex atlases [28]. The expert outputs are subsequently aligned with frozen CLIP embeddings, more details in Appendix A.
@@ 2.4 Dynamic Conditioning via Time and Space Routing
Generating images via diffusion is an iterative process. We leverage the hierarchical and specialized nature of the MoE embeddings dynamically, which allows the model to potentially focus on different levels of abstraction (semantic vs. detail) processed by different specialized networks at different stages of generation, mirroring coarse-to-fine processing theories [29]. We introduce a time and space routing mechanism (Figure 2b) to integrate the MoE expert embeddings ($E$) into the SDXL U-Net ($\epsilon_{\theta}$).
@@ 3.1 Settings and Implementations
We evaluate MoRE-Brain using the Natural Scenes Dataset (NSD) [30], a large-scale fMRI dataset comprising brain responses from 8 subjects viewing over 30,000 distinct natural images. We adhere to the standardized training and testing splits proposed in prior work [6]. Further details on NSD and our fMRI data preprocessing pipeline are available in Appendix B.
We compare MoRE-Brain with two methods: the recent state-of-the-art MindEye2 [6] and MindBridge [31]. To assess the extent to which models leverage fMRI information versus relying on generative priors, we adopt the bottleneck analysis proposed by [8], evaluating performance across varying bottleneck sizes imposed on the fMRI-derived features. We include high-level metrics like DreamSim [32] and InceptionV3 [6, 33], and a low-level metric SSIM. We also measure the cosine similarity between decoded embeddings and the ground truth CLIP embedding as in [8], thereby isolating the decoder’s performance from the generative model. For each method, we report a "reconstruction ceiling" (using ground truth image latents instead of fMRI-predicted latents) and a "random baseline" (using random fMRI as input). More details on implementation are in Appendix A.
Figure 4: Quantitative performance across varying bottleneck sizes. MoRE-Brain shows a consistent and notable performance drop across most metrics as bottleneck size decreases, indicating strong reliance on fMRI information. However, MindEye2 and MindBridge exhibit less degradation, particularly MindBridge on SSIM and InceptionV3, suggesting a greater influence of learned priors. Notably, MindEye2 maintains high CLIP-Cos even at extreme bottlenecks, while others show expected decline with information restriction.
@@
First, we evaluate the overall capabilities of fMRI-to-image reconstruction. Figure 3 presents qualitative examples of reconstructions for Subject 1, where MoRE-Brain faithfully recovers visual stimuli. Quantitatively, Figure 4 shows average performance across all subjects. MoRE-Brain’s performance is competitive with state-of-the-art methods. We also show reconstructions from single level of expert embeddings, see Appendix E.
@@ 3.3 Cross-Subject Generalization with Limited Data
Beyond robust fMRI-driven reconstruction, MoRE-Brain’s architecture, particularly its separation of shared experts from subject-specific routers, allows for efficient generalization with limited data. To demonstrate this, we first train the full model from scratch on subject 1, then freeze the weights of all experts and fine-tune only the fMRI routers (Eq. 2) with 2.5% (1 session), 10% (4 sessions), 25% (10 sessions), 50% (20 sessions), and all new subjects’ data (40 sessions). See results in Appendix C.
@@
[…]
@@ 3.5.1 Overall Brain Contributions to Reconstruction
To identify brain regions crucial for reconstruction across subjects, we compute voxel attributions using GradientSHAP [34] for the entire MoRE-Brain. The attribution maps (1000 per subject, 8000 total) are then decomposed using Independent Component Analysis (ICA) [35] into 32 spatial ICs, representing distinct co-varying patterns of functional activity contributing to the decoding.
Furthermore, by correlating IC activations (derived from the ICA mixing matrix) with the semantic content of the input stimuli (12 COCO supercategories [36]), we find that specific ICs exhibit significant preferences for object categories (Figure 7). For example, ICs strongly correlated with "appliance" consistently involve parieto-occipital regions. (Details on IC labeling with Schaefer atlas [37] in Appendix G).

### v3

@@ Abstract
Decoding visual experiences from fMRI offers a powerful avenue to understand human perception and develop advanced brain-computer interfaces. However, current progress often prioritizes maximizing reconstruction fidelity while overlooking interpretability, an essential aspect for deriving neuroscientific insight. To address this gap, we propose MoRE-Brain, a neuro-inspired framework designed for high-fidelity, adaptable, and interpretable visual reconstruction. MoRE-Brain uniquely employs a hierarchical Mixture-of-Experts architecture where distinct experts process fMRI signals from functionally related voxel groups, mimicking specialized brain networks. The experts are first trained to encode fMRI into the frozen CLIP space. A finetuned diffusion model then synthesizes images, guided by expert outputs through a novel dual-stage routing mechanism that dynamically weighs expert contributions across the diffusion process. MoRE-Brain offers three main advancements: First, it introduces a novel Mixture-of-Experts architecture grounded in brain network principles for neuro-decoding. Second, it achieves efficient cross-subject generalization by sharing core expert networks while adapting only subject-specific routers. Third, it provides enhanced mechanistic insight, as the explicit routing reveals precisely how different modeled brain regions shape the semantic and spatial attributes of the reconstructed image. Extensive experiments validate MoRE-Brain’s high reconstruction fidelity, with bottleneck analyses further demonstrating its effective utilization of fMRI signals, distinguishing genuine neural decoding from over-reliance on generative priors. Consequently, MoRE-Brain marks a substantial advance towards more generalizable and interpretable fMRI-based visual decoding. Codes are publicly available: https://github.com/yuxiangwei0808/MoRE-Brain.
@@ 1 Introduction
Understanding brain visual systems and human perception has been a central interest in the field of neuroscience wei2025symptomatic ; wei2025hierarchical ; xiao2025describe . The human visual system transforms perceived scenes into rich, meaningful perception through a complex cascade of hierarchical and specialized processing ganis2004brain . With the advancement in the diffusion models podell2023sdxl and representational space like CLIP radford2021learning or visual prompt tuning xiao2025visual , numerous studies propose to decode visual stimuli from non-invasive brain signals such as functional MRI (fMRI) and EEG, aiming to unravel the complex mechanism under perception and advance the practical brain-computer interface. These methods usually leverage strong priors from existing generative models and train dedicated decoders to decode brain signals to frozen representational space, reconstructing natural scenes with impressive structural accuracy scotti2024mindeye2 ; luo2023brain ; mayo2024brainbits ; zhou2024clip ; lu2023minddiffuser ; diao2025ft2tf .
@@
Despite this progress, a significant gap persists between technical feats and neuroscientific insight. Current leading methods often treat fMRI signals as relatively monolithic inputs, lacking architectural designs that reflect established principles of neural processing. For instance, several works manually segment the brain into broad regions like the higher visual and lower visual cortex, then decode semantic and low-level spatial information from them separately fang2020reconstructing ; takagi2023high ; joo2024brain ; diao2025temporal . However, this simplification overlooks the intricate nature of the visual pathway, which involves multiple hierarchical processing stages across numerous specialized cortical areas, each handling different facets of visual information perry2014feature ; brewer2002visual . Furthermore, leading research commonly encounters two critical bottlenecks: cross-subject generalization and limited mechanistic interpretability. Due to inter-individual brain variability, models often require subject-specific training. While approaches like mapping subject data to a shared space scotti2024mindeye2 ; ferrante2024through ; scotti2023reconstructing or utilizing subject-specific tokens zhou2024clip exist, they typically necessitate extra training and are difficult to generalize to new subjects. More fundamentally, the predominant focus on maximizing decoding performance often obscures how different neural computations contribute to the final reconstruction. This lack of transparency hinders our ability to validate these models against neuroscientific knowledge or use them to deepen our understanding of the visual system itself.
@@
To bridge the gap, we introduce MoRE-Brain (Mixture-of-Experts Routed Brain decoder), a novel visual decoding framework that is explicitly designed to mirror the hierarchical and specialized nature of the human visual pathway felleman1991distributed to overcome key limitations in generalization and interpretability inherent in less biologically constrained approaches. MoRE-Brain builds upon a hierarchical Mixture-of-Experts (MoE) architecture to first map fMRI activity patterns to image and text priors. Instead of relying on a single, monolithic network, MoRE-Brain employs multiple layers of "expert" subnetworks. Within each layer, trainable "router" networks learn to assign specific brain regions (voxel groups) to the experts best suited to process their signals. This allows each expert to specialize in decoding information from distinct neural populations, analogous to functional specialization in the cortex. Addressing the challenge of cross-subject generalization, we draw from findings suggesting that significant inter-individual variation arises from spatial differences in functional network topography rather than fundamental computational differences anderson2021heritability ; savage2024dissecting . Based on this, we hypothesize that the core visual decoding computations (the experts) can be shared across subjects, while subject-specific adjustments are primarily needed for the routers that map individual brain topographies to these experts. Our experiments validate this, demonstrating that by freezing expert weights and fine-tuning only the routers, MoRE-Brain achieves strong cross-subject performance with minimal subject-specific data, offering a more scalable approach than retraining entire models.
@@
Furthermore, effectively utilizing the disentangled information captured by these specialized experts within the iterative image generation process of diffusion models presents another challenge. This mechanism operates dynamically across both time and space: (1) A Time Router selects which hierarchical level of expert embeddings to prioritize at different diffusion timesteps. This potentially reflects the coarse-to-fine dynamics of visual processing, where global scene layout might be established before finer details emerge skyberg2022coarse . (2) A Space Router then dynamically modulates the spatial influence of the outputs from individual experts within the selected level. This mimics how the brain might integrate features processed by different specialized areas (e.g., form, color, motion) into a spatially coherent percept goodale1992separate . Critically, the dual-routing system not only guides image generation but also provides a mechanistic lens into the reconstructing process. By observing which experts and levels are routed at different times and for images with different semantic concepts, we gain insight into how distinct, modeled neural sources dynamically contribute to the final reconstruction.
@@ 2.1 Preliminaries: Latent Diffusion Models and Conditioning
Our reconstruction module leverages latent diffusion models (LDMs), specifically Stable Diffusion XL (SDXL) podell2023sdxl , which are generative models operating in a compressed latent space ho2020denoising ; rombach2022high . LDMs learn to reverse a diffusion process that gradually adds Gaussian noise to latent representations ($z_{0}$) over discrete timesteps $t=1,\dots,T$. The forward process is defined as $z_{t}=\sqrt{\bar{\alpha}_{t}}z_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$, where $\bar{\alpha}_{t}$ is a noise schedule and $\epsilon\sim\mathcal{N}(0,\mathbf{I})$. The core of the LDM is a denoising network $\epsilon_{\theta}$ (typically a U-Net ronneberger2015u ), trained to predict the noise $\epsilon$ added at timestep $t$, given the noisy latent $z_{t}$ and conditioning information $c$:
[…]
Conditioning information $c$ is typically derived from various modalities. For text-to-image synthesis, $c$ often consists of embeddings $\tau_{text}(c_{text})$ from a pre-trained text encoder (e.g., CLIP radford2021learning ). Image conditioning can be achieved using IP-Adapter ye2023ip , where features $\tau_{img}(c_{img})$ are extracted from a reference image via a pre-trained image encoder (e.g., CLIP image encoder) and mapped through an adapter network. Both $\tau_{text}(c_{text})$ and $\tau_{img}(c_{img})$ are integrated into $\epsilon_{\theta}$ usually via cross-attention mechanisms. In MoRE-Brain, we adapt this conditioning mechanism to use fMRI-derived embeddings as image and text priors.
@@ 2.2 MoRE-Brain Framework Overview
  • Dynamically Conditioned Image Generation: We introduce a novel Time and Space routing mechanism to dynamically select and integrate the multi-level expert embeddings produced in Stage 1 to guide the diffusion model’s (SDXL’s) denoising process at each timestep $t$. During this stage, the MoE fMRI encoder (trained in Stage 1) is frozen, while the SDXL U-Net (via LoRA hu2022lora ) and the Time and Space routers are fine-tuned.
@@ 2.3 Hierarchical MoE for fMRI Encoding
For progressively finer specialization, outputs $O_{j}^{(l)}$ from experts at level $l$ serve as input $X^{(l+1)}$ to the routers at the next level $(l+1)$. We start with $e_{0}=2$ experts at the first level, and each expert’s output path is routed to 2 new experts at the subsequent level. We use a total of $L=4$ levels, resulting in $2^{4}=16$ experts at the final level. This number was empirically chosen to approximate the number of distinct functional ROIs identified in recent visual cortex atlases rosenke2021probabilistic . The expert outputs are subsequently aligned with frozen CLIP embeddings, more details in Appendix A.
@@ 2.4 Dynamic Conditioning via Time and Space Routing
Generating images via diffusion is an iterative process. We leverage the hierarchical and specialized nature of the MoE embeddings dynamically, which allows the model to potentially focus on different levels of abstraction (semantic vs. detail) processed by different specialized networks at different stages of generation, mirroring coarse-to-fine processing theories goffaux2011coarse . We introduce a time and space routing mechanism (Figure 2b) to integrate the MoE expert embeddings ($E$) into the SDXL U-Net ($\epsilon_{\theta}$).
@@ 3.1 Settings and Implementations
We evaluate MoRE-Brain using the Natural Scenes Dataset (NSD) allen2022massive , a large-scale fMRI dataset comprising brain responses from 8 subjects viewing over 30,000 distinct natural images. We adhere to the standardized training and testing splits proposed in prior work scotti2024mindeye2 . Further details on NSD and our fMRI data preprocessing pipeline are available in Appendix B.
We compare MoRE-Brain with two methods: the recent state-of-the-art MindEye2 scotti2024mindeye2 and MindBridge wang2024mindbridge . To assess the extent to which models leverage fMRI information versus relying on generative priors, we adopt the bottleneck analysis proposed by mayo2024brainbits , evaluating performance across varying bottleneck sizes imposed on the fMRI-derived features. We include high-level metrics like DreamSim fu2023dreamsim and InceptionV3 scotti2024mindeye2 ; szegedy2016rethinking , and a low-level metric SSIM. We also measure the cosine similarity between decoded embeddings and the ground truth CLIP embedding as in mayo2024brainbits , thereby isolating the decoder’s performance from the generative model. For each method, we report a "reconstruction ceiling" (using ground truth image latents instead of fMRI-predicted latents) and a "random baseline" (using random fMRI as input). More details on implementation in Appendix A.
Figure 4: Quantitative performance across varying bottleneck sizes. MoRE-Brain shows a consistent and notable performance drop across most metrics as the bottleneck size decreases. Moreover, its performance witnesses the most significant drop after adding a bottleneck, indicating that the bottleneck is particularly detrimental to the learned rich information and the model heavily relies on the fMRI data. However, MindEye2 and MindBridge exhibit less degradation, particularly MindBridge on SSIM and InceptionV3, suggesting a greater influence of learned priors. Notably, MindEye2 maintains high CLIP-Cos even at extreme bottlenecks, while others show expected decline with information restriction.
@@
First, we evaluate the overall capabilities of fMRI-to-image reconstruction. Figure 3 presents qualitative examples of reconstructions for Subject 1, where MoRE-Brain faithfully recovers visual stimuli. Quantitatively, Figure 4 shows average performance across all subjects. MoRE-Brain’s performance is competitive with state-of-the-art methods. We also show reconstructions from a single level of expert embeddings, see Appendix E.
@@ 3.3 Cross-Subject Generalization with Limited Data
MoRE-Brain’s separation of shared experts from subject-specific routers allows for efficient generalization with limited data. After training the model on subject 1, we freeze the model’s experts and only finetune the routers to generalize to other subjects. This massively reduces the parameters during finetuning. We compare the total parameters of MoRE-Brain and the two baselines, as in Table 1. Note that for MindBridge, we applied its ’reset-tuning’ strategy.
Table 1: Compare the total number of parameters and trainable parameters during finetuning.
Total Params (M) Trainable (%)
MindEve2 729.3 100
MindBridge 552.9 98.48
MoRE-Brain 293.4 44.84
To further demonstrate MoRE-Brain’s generalizing performance with limited data, we first train the full model from scratch on subject 1, then freeze the weights of all experts and fine-tune only the fMRI routers (Eq. 2) with 2.5% (1 session), 10% (4 sessions), 25% (10 sessions), 50% (20 sessions), and all new subjects’ data (40 sessions). See results in Appendix C.
@@
$\mathcal{T}$ $\mathcal{T}$ (TS) $\mathcal{I}$ $\mathcal{I}$ (TS) $\mathcal{T}$$\mathcal{I}$ $\mathcal{T}$$\mathcal{I}$ (S) $\mathcal{T}$$\mathcal{I}$ (T, sched) $\mathcal{T}$$\mathcal{I}$ (T, hard) $\mathcal{T}$$\mathcal{I}$ (T, soft) $\mathcal{T}$$\mathcal{I}$ (TS)
@@ 3.5.1 Overall Brain Contributions to Reconstruction
To identify brain regions crucial for reconstruction across subjects, we compute voxel attributions using GradientSHAP lundberg2017unified for the entire MoRE-Brain. The attribution maps (1000 per subject, 8000 total) are then decomposed using Independent Component Analysis (ICA) hyvarinen2000independent into 32 spatial ICs, representing distinct co-varying patterns of functional activity contributing to the decoding.
Furthermore, by correlating IC activations (derived from the ICA mixing matrix) with the semantic content of the input stimuli (12 COCO supercategories lin2014microsoft ), we find that specific ICs exhibit significant preferences for object categories (Figure 7). For example, ICs strongly correlated with "appliance" consistently involve parieto-occipital regions. (Details on IC labeling with Schaefer atlas schaefer2018local in Appendix G).